# Quantum-Informed Machine Learning for Chaotic Systems

### v2

@@ II.3 Turbulent Channel Flow Inflow Generation
Collectively, these results demonstrate that embedding Q-priors significantly improves long-term prediction results and statistical alignment of classic machine learning models in chaotic systems. Even with modest quantum resources, that is no more than 15 qubits and under 300 parameters (specifically shown in method F, Supplementary Information S2.2 and S2.3), our framework captures essential dynamics of high-dimensional flows and enables practical hybrid learning pipelines for scientific applications. By anchoring predictions to Q-prior, the QIML model improves statistical fidelity and stability, laying the groundwork for practical quantum–classical synergies in the modelling of turbulent and other chaotic physical systems.
III Discussion
This work introduces a quantum-informed machine learning framework for learning chaotic dynamical systems. In QIML, a QCBM is trained on superconducting hardware to learn invariant statistical properties from high-dimensional flow data, and the resulting Q-prior is integrated with a classical transformer-based machine learning model. The QIML framework is validated across three representative chaotic systems, demonstrating consistent improvements in long-term stability, physical interpretability, and statistical fidelity compared to classical baselines without Q-prior. The consistent improvements across qualitatively distinct systems suggest that the proposed architecture generalises well and may be adapted to a broader class of nonlinear dynamical systems.
A central contribution of this work lies in its demonstration of parameter-efficient learning. Using fewer than 300 trainable parameters, the QCBM captures statistical distributions over a state space with more than 30,000 dimensions. This parameter economy arises because the QCBM explores a high-dimensional Hilbert space, allowing it to learn long-range velocity correlations with only a few hundred rotation angles. Crucially, the model is trained with a sample-based MMD loss, so it can infer these correlations directly from finite trajectories without requiring an explicit form of the underlying, often unknown, probability density. The proposed QIML architecture establishes a scalable and hardware-compatible pathway for embedding Q-priors into classical machine learning pipelines. By separating the learning of invariant flow statistics from the time-stepping dynamics, and compressing these complex statistics into highly memory-efficient Q-priors, the framework improves both generalisation and training stability without requiring full end-to-end optimisation. Nonetheless, limitations remain. To reduce the impact of hardware noise and enable stable training, the QCBM is optimised separately and remains fixed during the training of the classical model. This architectural constraint arises from the limited scalability and noises of current superconducting devices, which make full end-to-end quantum training infeasible at present. Although integrated as a static distributional prior, the QCBM captures complex statistical structures, potentially including entanglement, that would require a large amount of classical parameters to approximate. While current QCBMs are limited by qubit count, circuit depth, and hardware noise, deeper architectures with structured entanglement (e.g. hardware-efficient ansätze) may further improve expressivity for highly anisotropic distributions.
@@ III Discussion
While this work demonstrates improvements in learning from digitally generated chaotic system data, we acknowledge the inherent limitations of digital computing in fully capturing the true continuum behaviour of chaotic systems [5, 6, 7, 8]. Our ground truth here is, by definition, a 32 bit floating-point approximation. Future work will explore the use of analogue computation to provide a more direct comparison with the true invariant measures, thus offering a more rigorous validation of our approach against the ultimate ground truth.
@@
QIML demonstrates that even current NISQ devices can contribute meaningfully to the modelling of high-dimensional dynamical systems. Without resorting to fault-tolerant hardware, lightweight quantum modules enhance classical surrogates by providing data-driven Q-priors that efficiently capture complex statistical structures and non-local correlations—features that are otherwise challenging for classical methods to acquire with comparable memory efficiency. The resulting gains in statistical fidelity and long-horizon stability, coupled with a substantial reduction in data storage requirements, point to an immediate and practical role for near-term quantum resources in scientific machine learning. The compact 10- to 15-qubit QCBM presented here exemplifies how such modules can be effectively integrated today. Looking forward, the hybrid architecture introduced here serves as an intermediate but essential step toward fully realising quantum machine learning pipelines. While our current framework integrates quantum priors into classical surrogate models, we envision future extensions in which quantum circuits not only generate priors but also perform end-to-end learning and generation. Given the exponentially larger Hilbert space accessible to quantum systems, such models are expected to possess superior expressivity for learning high-dimensional generative tasks. As quantum hardware continues to advance, in terms of qubit count, coherence time, and error mitigation, the prospect of replacing classical generative backbones with fully quantum circuits becomes increasingly viable. Thus, the parameter and memory efficiency demonstrated in this work suggest a pathway for quantum modules to ultimately replace entire sections of large classical models.

### v3

@@ III.4 Turbulent channel flow inflow generation
Collectively, these results demonstrate that embedding Q-Priors significantly improves long-term prediction results and statistical alignment of classic machine learning models in high-dimensional dynamical systems. Even with modest quantum resources, that is, no more than 15 qubits and under 300 parameters (specifically shown in method F, Appendix B), our framework captures essential dynamics of high-dimensional flows and enables practical hybrid learning pipelines for scientific applications. By anchoring predictions to Q-Prior, the QIML model improves statistical fidelity and stability, laying the groundwork for practical quantum–classical synergies in the modeling of turbulent and other chaotic physical systems.
IV Comparative Analysis of QIML against Machine Learning Baselines
To rigorously quantify the contribution of our quantum-informed machine learning framework, we conduct a comparative analysis of our full QIML architecture against key baseline models. The aim is to show that the stability and statistical fidelity observed in our results arise from the Q-Prior guidance, rather than being solely attributable to the autoregressive classical ML model’s architecture. Both the QIML framework and the classical baselines are trained on exactly the same raw dataset. The fundamental difference lies in the architectural approach instead of unequal access to data. Here, we consider two primary classes of baselines: (i) a classical state-of-the-art Koopman-based autoregressive machine learning model trained without the Q-Prior, and (ii) neural operators for nonlinear PDEs and chaotic systems such as the Fourier neural operator and the Markov neural operator [78].
As demonstrated in Fig. 7, our classical model without the Q-Prior consistently fails to maintain long-term stability. While capable of accurate short-term predictions, it inevitably succumbs to error accumulation, leading to prediction divergence and a failure to reproduce the system’s invariant statistical measures. The comparison with neural operators provides a more stringent test. FNOs and MNOs are known for their strong performance in learning resolution-independent solutions to nonlinear PDEs and chaotic systems. We trained both models on the same datasets used for our framework. We observed that for short-term predictions, their accuracy is competitive. However, when tasked with long-horizon autoregressive rollouts, the classical Koopman-based machine learning model, the FNO and MNO models exhibit the same fundamental weakness: a gradual divergence from the ground truth and a failure to preserve the correct spatio-temporal evolution of the chaotic system. This failure is particularly striking given their substantial size; both the FNO and MNO models employ an order of magnitude more trainable parameters than our QIML framework, as detailed in the comparison in the Table 3.
This comparative analysis leads to a clear conclusion: the key enabling component of our framework is the Q-Prior integrated into the machine learning. The long-term instability is a fundamental challenge for even powerful classical architectures like FNOs and MNOs. By providing a physically grounded, memory-efficient statistical regularizer, the Q-Prior confers the necessary stability and long-term fidelity that are otherwise absent. This result underscores the unique and practical advantage of our QIML architecture.
V Q-Prior memory efficiency
A defining feature of the proposed framework is its capacity to extract high-dimensional physical information using significantly fewer trainable parameters than conventional deep learning models. The Q-Prior, operating on only 10–15 qubits and comprising fewer than 300 trainable parameters, successfully captures invariant distributions over spatial domains with dimensionality exceeding . In contrast, a classical neural network capable of learning an equally expressive distribution over the same domain would require at least parameters, assuming a minimal fully connected architecture with comparable granularity. For example, on a grid of spatial points, the lightest fully-connected (FC) network that can approximate an arbitrary probability map with standard universal-approximation accuracy must allocate weights [79, 80]. Even with an aggressively thinned hidden layer, this still amounts to tunable parameters—three orders of magnitude more than the parameters used by our ten-qubit Q-Prior (see Appendix B). Moreover, classical models lack access to quantum entanglement, limiting their ability to compactly encode non-local correlations and higher-order statistical dependencies prevalent in turbulent or chaotic systems.
However, possessing a parameter-efficient quantum representation is only the first step; leveraging it to guide a classical machine learning model is a distinct and significant challenge. We address this by designing a composite loss function that seamlessly integrates the statistical information from the Q-Prior into the classical model’s training process. This enables a true “quantum-informed” learning paradigm where the quantum component actively regularizes the classical dynamics, leading to superior results as demonstrated in this work. As shown in the Table 1, the successful implementation of this hybrid integration unlocks a further, highly practical potential of quantum computing: memory efficiency in data storage and transfer.
Specifically, for the turbulent–channel dataset, retaining all mid-plane snapshots in double precision requires roughly . Archiving the corresponding Q-Prior checkpoint for each snapshot ( per file) occupies only , a compression greater than two orders of magnitude while preserving the relevant flow statistics. Comparable savings are obtained for Kolmogorov flow ( parameters, per file, in total) and Kuramoto–Sivashinsky data ( parameters, per file, in total). Full storage calculations are shown and reported in Appendix H. Such dramatic savings make the Q-Prior attractive whenever data are expensive to generate or archive, while still providing a statistically faithful, low-parameter regularizer for the classical Koopman machine learning.
Crucially, this provides a blueprint for future hybrid computing infrastructures where the quantum processors might be co-located with every classical computer. Instead of transferring voluminous raw datasets, which can amount to hundreds of gigabytes or more, one would only need to transfer the compact Q-Prior. As our results show, this corresponds to a data storage reduction of over two orders of magnitude. Furthermore, this vision is made practical by a key finding that directly addresses the number of measurement shots, a widely scrutinized issue for NISQ-era algorithms. While the results presented in this work are based on a high shot count (20,000 per iteration) to demonstrate the framework’s maximum potential, we found that the classical model does not require a perfectly resolved, exponentially large distribution to be effective. In fact, reducing the measurement count to 8,000 shots had a negligible impact (less than 5%) on the final model’s performance, showing that an approximate prior is sufficient to act as a powerful guiding signal. Crucially, the Q-Prior is trained only once in an offline stage, rather than being re-evaluated during every rollout or trained iteratively within the ML framework like QEML. This one-time training greatly reduces the overall computational burden and differentiates our approach from iterative hybrid quantum–classical schemes that require repeated quantum calls throughout the learning process. Thus, QIML establishes a viable and economically attractive framework: a central, high-fidelity quantum processor performs the intensive, one-time training of a master Q-Prior, whose compact parameters are then distributed at low cost to remote systems. This paradigm may serve as a template for future cloud quantum platforms, where the quantum processors provide powerful statistical guidance and data support for large-scale classical models.
Dynamical System | Raw data | Q-Prior | Compression |
(full set) | (full set) | (ratio) | |
Kuramoto–Sivashinsky | |||
2D Kolmogorov flow | |||
Turbulent channel flow |
VI On the Nature of Quantum “Advantage” in QIML
The term “quantum advantage” must be approached with caution, particularly in the context of quantum computing and quantum machine learning [65, 81]. As recent analyses have highlighted, many claims of quantum advantage are later matched or surpassed by sophisticated classical algorithms, revealing that the initial gap was due to the immaturity of classical methods rather than a fundamental quantum superiority [65]. We have deliberately refrained from prominently using this term because of its ambiguity. With the exception of algorithms like Shor’s [82], which leverage the fundamental nature of quantum mechanics in a provably superior way, many claims of advantage are eventually challenged. Even powerful algorithms such as the Harrow–Hassidim–Lloyd (HHL) algorithm have faced scrutiny [83], with new classical works emerging that achieve comparable performance in specific contexts [65]. Here, we do not assert that QIML holds an unassailable advantage. Instead, we posit that its effectiveness stems from a foundational quantum principle that, in this application, has not been classically replicated.
The foundational advantage exploited by our Q-Prior stems from the ability of an -qubit quantum state to represent a classical vector in a dimensional space, which is considered as a genuine quantum advantage [67]. This provides an exponential advantage in information storage capacity. This type of advantage has found recent applications in quantum machine learning, such as for computing quadratic functions of data [84]. However, this potential is famously constrained by Holevo’s bound, which dictates that only classical bits of information can be reliably extracted from an n-qubit state through measurement [85]. If full reconstruction of the classical information were required, the need for an exponential number of measurements would negate the storage advantage entirely.
In ergodic theory, the invariant measure governs the long-time statistical behavior of chaotic dynamics, even though individual trajectories diverge exponentially [86]. The central insight of the QIML framework is to circumvent Holevo’s bound by extracting only this essential invariant measure without reconstructing the full dataset, yielding a compressed yet sufficient representation of the system. This low-dimensional representation is precisely what is needed to regularize the classical ML, guiding it away from non-physical trajectories without needing to know every detail of the system’s state. Therefore, the advantage demonstrated here is a representational and memory advantage that translates directly into a performance advantage.
Memory Efficiency: The Q-Prior leverages quantum entanglement to capture complex, non-local correlations within the invariant measure using a few parameters. This translates to a dramatic, orders-of-magnitude reduction in data storage, as quantified in Table 1.
Performance Gain: This compact and physically rich representation provides the crucial regularization that leads to long-horizon stability and statistical fidelity, a task at which larger, more complex classical models like FNO and MNO demonstrably fail for this class of problems.
While it is an open question whether other classical models could be adapted to leverage the Q-Prior as effectively, our results suggest that the co-design of the hybrid architecture is key. We anticipate that our results may inspire new classical approaches. For instance, using large generative models to learn the invariant measure, that could achieve comparable predictive performance. However, we argue that such methods are unlikely to replicate the foundational source of the advantage demonstrated here: the physical ability to logarithmically compress high-dimensional data and extract the necessary guiding information without an exponential measurement cost. This is possible because the invariant measure appears to be a sufficiently rich representation of the essential physics of chaos. Capturing its key features provides the necessary guidance to unlock the observed gains in performance and, crucially, in memory efficiency. This efficiency is a direct, physical consequence of the quantum representation, and while our theoretical understanding of the measure’s precise role is still developing, the empirical results provide compelling evidence for this paradigm’s value.
Another notable point is that much of that research begins with a search for theoretical advantage, often struggling to connect with practical, “useful” applications [65]. In contrast, our investigation was born from a practical necessity: to solve the long-horizon prediction problem for chaotic systems, where leading classical methods fail. By showing that a 10-qubit quantum device can provide a unique resource, this work provides a compelling example of a useful quantum “advantage” and motivates the further development of both quantum and classical algorithms for scientific machine learning.
VII Immediate and Future Impact of QIML
@@
To reduce the impact of hardware noise and enable stable training, the quantum generator is optimized separately and remains fixed during the training of the classical model. This architectural constraint arises from the limited scalability and noise of NISQ superconducting devices, which make full end-to-end quantum training infeasible at present. For near-term work, further improvements are possible within this offline training paradigm. While current Q-Priors are limited by qubit count, circuit depth, and hardware noise, deeper architectures with structured entanglement (e.g. hardware-efficient ansätze) may further improve expressivity for highly anisotropic distributions. Besides, while this work demonstrates improvements in learning from digitally generated high-dimensional dynamical system data, we acknowledge the inherent limitations of the digital computing part of our method in fully capturing the true continuum behavior of high-dimensional dynamical systems [5, 6, 7, 8]. Our ground truth here is, by definition, a 64-bit floating-point approximation. Future work will explore the use of analogue computation to provide a more direct comparison with the true invariant measures, thus offering a more rigorous validation of our approach against the ultimate ground truth.
@@
Although there are limitations from NISQ hardware, in contrast to prevailing scepticism about the immediate applicability of quantum learning algorithms [36, 81], QIML demonstrates that even current quantum devices can contribute meaningfully to the modeling of high-dimensional dynamical systems. Without resorting to fault-tolerant hardware, lightweight NISQ-based quantum modules enhance classical machine learning models by providing data-driven Q-Priors that efficiently capture complex statistical structures and non-local correlations—features that are otherwise challenging for classical methods to acquire with comparable memory efficiency. The resulting gains in statistical fidelity and long-horizon stability, coupled with a substantial reduction in data storage requirements, point to an immediate and practical role for near-term quantum resources in scientific machine learning. The compact 10- to 15-qubit Q-Prior presented here exemplifies how such modules can be effectively integrated today.
Looking forward, the QIML architecture introduced here serves as an intermediate but essential step toward fully realizing quantum machine learning pipelines. While our current framework integrates quantum priors into classical machine learning models, we envision future extensions in which quantum circuits not only generate priors but also perform end-to-end learning and generation for high-dimensional datasets. Given the exponentially larger Hilbert space accessible to quantum systems, such models are expected to possess superior expressivity for learning high-dimensional generative tasks.
Furthermore, this work opens up more exciting avenues for future research. A key question is how the framework performs with sparse or partial datasets, a common scenario in real-world applications where obtaining a complete ground-truth solution is impossible. Our current results are already promising in this regard, as the Q-Priors were successfully trained using only a subset of the available data (see section III) while still providing effective guidance. This suggests the potential for a more profound, out-of-distribution generalization [87]. For instance, future work could explore whether a single Q-Prior, trained on the invariant measure of one system, could successfully regularize a model for a different system that shares similar statistical properties but has different underlying dynamics. The successful demonstration of such a “universal” prior would have significant implications, potentially allowing one Q-Prior to guide multiple machine learning models, further enhancing the practical value and efficiency of the QIML paradigm.